OpenStudy (dallaslifebaby01):

If f(x) = 3x + 7, what is f(-2)? -2 13 1 -1

OpenStudy (caozeyuan):

f(-2)=3*(-2)+7=-6+7=1

OpenStudy (dallaslifebaby01):

thanks